OpenStudy (anonymous):

integrate from 0 to pi/2 (sinx + cosx)/(9+16sin2x) dx

OpenStudy (anonymous):

Recall sin 2x= 2[[(cos)^2]x]-1

OpenStudy (anonymous):

yup.. what nxt?

OpenStudy (anonymous):

Choose your u.

OpenStudy (anonymous):

can you solve it out buddy.. i seem to get nowhere!

OpenStudy (anonymous):

You have to make an attempt. Don't get overwhelm. Your task is simple: choose a u. You could be right, you could be wrong. You won't harm anything.

OpenStudy (anonymous):

\[\sin x+\cos x \over 9+16(2\cos ^{2}x-1)\] \[\sin x+ \cos x \over 32\cos ^{2}x-7\] what do you think i should select as u

OpenStudy (anonymous):

We're not getting anywhere. I asked you twice: choose a u.

OpenStudy (anonymous):

i have no idea buddy.. i ll try later

OpenStudy (watchmath):

I am curious how we can really go from there. BTW \(\sin(2x)\neq 2\cos^2(x)-1\). But \(\cos(2x)=2\cos^2(x)-1\)

OpenStudy (anonymous):

lol... i didnt evn realize that! rats!

OpenStudy (anonymous):

Good catch, watchmath. Apparently this question appears frequently in India exams. Solution depends on the relation 1 - sin 2x=(sinx - cos x)^2. Let sin x - cos x = t (cos x + sin x) dx = dt t^2= (sin x - cos x)^2 = 1 - sin 2x sin 2x = 1-t^2.\[\int\limits_{0}^{\pi/2}(\sin x + \cos x)/(9+16 \sin 2x)\]=\[\int\limits_{0}^{\pi/2}dt/(9+16(1-t ^{2})\]